# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v1

@@ 4.3. $\mathbb{Z}$-valued Solutions and Finite-Difference Equations
where $g_{\mathbb{Z}}$ denotes the left-inverse of the function $u\mapsto u+f_{\mathbb{Z}}(u)$. For such a law $\mathbf{P}$, all the assumptions of Theorem 1 are satisfied. In particular, as long as $(f_{+},f_{-})$ is not identically zero, either $n^{-\frac{1}{2}}X^{(n)}$ or $n^{-\frac{1}{3}}X^{(n)}$ converges to a nondegenerate limit.
@@ 4.4. Hipster Random Walks and Cooperative Motion
In fact, by this reasoning, it is also possible to draw a connection with the other RDE considered in [1], called the totally asymmetric hipster random walk. This is the RDE in $\mathbb{Z}$ with the one-step update determined by two independent Bernoulli random variables $\Theta,\Xi$ via the rule
@@
In view of the proposition above, the $\text{Beta}(2,2)$ limit theorems for the symmetric RDE’s considered in [1] and [3] follow as a special case of this paper’s main theorem (Theorem 1) and Corollary 1. The $\text{Beta}(2,1)$ limit theorems for asymmetric RDE’s obtained in [1] and [2] also follow, albeit only in the parameter regimes identified above. Interestingly, in the asymmetric case, those works obtain $\text{Beta}(2,1)$ limits, even in the nonmonotone regime.
@@ 5. Scaling Limit
This section proves the main result, Theorem 1. In fact, a slightly more general version is presented here, which allows for extended-real-valued solutions and diffuse initial conditions. In order to state the theorem efficiently, as in the introduction, let $\alpha=2$ if $\sigma\neq 0$ and $\alpha=3$, otherwise, and define the nonlinearity $\mathcal{F}$ appearing in the limiting PDE analogously:
(48) $\mathcal{F}(p,X)=\left\{\begin{array}[]{r l}\sigma|p|^{2},&\text{if}\,\,\sigma\neq 0,\\ a|p|X,&\text{otherwise.}\end{array}\right$.
The theorem that follows involves the notion of viscosity solutions of parabolic PDE’s. The relevant definitions and existence and uniqueness results are reviewed in Appendix A below.
@@
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds and fix an arbitrary $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$.
For each $\epsilon>0$, let $\{X_{\epsilon}^{(n)}\}$ be an extended-real-valued solution of the RDE (9) and let $\{F^{\epsilon}_{n}\}$ be the associated sequence of CDF’s. Let $F_{\epsilon}:\mathbb{R}\times[0,\infty)\to[0,1]$ be the rescaled CDF given by
$F_{\epsilon}(x,t)=F^{\epsilon}_{[\epsilon^{-\alpha}t]}(\epsilon^{-1}x).$
If convergence holds as $\epsilon\to 0$ at the initial time in the sense that
$F_{\epsilon}(x,0)\to F_{\text{in}}(x)\quad\text{for each point of continuity}\,\,x\,\,\text{of}\,\,F_{\text{in}},$
@@
$F_{\epsilon}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times(0,\infty),$
where $F$ is the unique bounded discontinuous viscosity solution of the initial-value problem
$\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
The theorem will be proved by following the approach of [6], suitably extended to monotone semigroups in the space $CDF(\overline{\mathbb{R}})$. Notice that if $T_{\epsilon}$ and $\mathscr{L}_{\epsilon}$ are the operators on $CDF(\overline{\mathbb{R}})$ defined by
(49) $(T_{\epsilon}G)(x)=(TG_{\epsilon})(\epsilon^{-1}x),\quad\text{where}\,\,G_{\epsilon}(x)=G(\epsilon^{-1}x),\quad\mathscr{L}_{\epsilon}=T_{\epsilon}-\text{Id},$
then the rescaled CDF $F_{\epsilon}$ above satisfies the equation
(50) $F_{\epsilon}(x,t+\epsilon^{\alpha})-F_{\epsilon}(x,t)=\mathscr{L}_{\epsilon}F_{\epsilon}(x,t).$
This motivates the consideration of one-parameter families of operators $\{T_{\epsilon}\}$ defined on $CDF(\overline{\mathbb{R}})$.
5.1. Abstract Framework
Let $\{T_{\epsilon}\}$ be a family of operators on $CDF(\overline{\mathbb{R}})$ with the following properties:
  • (i)
    Monotonicity: If $F\leq G$ pointwise in $\mathbb{R}$, then $T_{\epsilon}F\leq T_{\epsilon}G$ also holds.
  • (ii)
    Commutation with addition of constants: If $F\in CDF(\overline{\mathbb{R}})$ satisfies $c_{-}\leq F\leq c_{+}$ pointwise for some constants $c_{-},c_{+}\in[0,1]$, then $T_{\epsilon}(F+c)=T_{\epsilon}F+c$ for any $c\in[-c_{-},1-c_{+}]$.
  • (iii)
    Stationarity at infinity: For any $F\in CDF(\overline{\mathbb{R}})$, $T_{\epsilon}F(+\infty)=F(+\infty)$ and $T_{\epsilon}F(-\infty)=F(-\infty)$.
Motivated by (50), let $\{\mathscr{L}_{\epsilon}\}$ be the operators on $CDF(\overline{\mathbb{R}})$ defined by
$\mathscr{L}_{\epsilon}=T_{\epsilon}-\text{Id}.$
In addition to (i) and (ii), assume that there is an exponent $\alpha>0$ and a continuous function $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, which is increasing in the second variable, for which the following property holds:
  • (iv)
    Consistency: If $\Phi\in CDF(\overline{\mathbb{R}})$ is smooth with bounded, uniformly continuous second derivatives, then
    $\lim_{\epsilon\to 0}\epsilon^{-\alpha}\mathscr{L}_{\epsilon}\Phi(x)=\mathcal{F}(\partial_{x}\Phi(x),\partial_{x}^{2}\Phi(x)),$
    where convergence holds uniformly with respect to $x\in\mathbb{R}$.
Note that (iii) implies $Tc=c$ for any constant (function) $c\in[0,1]$, and, thus, by (iv),
(51) $\mathcal{F}(0,0)=0.$
Under these assumptions, the following convergence result holds for rescaled solutions $\{F^{\epsilon}_{n}\}$ of the recursion $F^{\epsilon}_{n+1}=T_{\epsilon}F^{\epsilon}_{n}$.
Theorem 3.
Assume that $\{T_{\epsilon}\}$ is a family of operators on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(iv) above. For each $\epsilon$, let $\{F^{\epsilon}_{n}\}$ be a sequence in $CDF(\overline{\mathbb{R}})$ generated by the recursion $F_{n+1}^{\epsilon}=T_{\epsilon}F_{n}^{\epsilon}$. If $F_{\epsilon}:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is the rescaled CDF defined by
$F_{\epsilon}(x,t)=F^{\epsilon}_{[\epsilon^{-\alpha}t]}(x),$
and if there is a continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ such that $F_{\epsilon}(x,0)\to F_{\text{in}}(x)$ pointwise as $\epsilon\to 0$, then
$F_{\epsilon}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times(0,\infty),$
where $F$ is the unique bounded continuous viscosity solution of the equation
(52) $\left\{\begin{array}[]{r l}\partial_{t}F=\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
Notice that the one-parameter family $\{T_{\epsilon}\}$ of interest in this paper (see (49)) satisfies assumptions (i)-(iii) according to Propositions 1 and 5. It only remains to verify condition (iv). That is the subject of the next proposition.
@@ 5.1. Abstract Framework
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, and consider the operator $\mathscr{L}$ given by (31). If $\Phi:\mathbb{R}\to[0,1]$ is a nondecreasing function with bounded, uniformly continuous second derivative, rescaled according to the rule $\Phi_{\delta}(x)=\Phi(\delta x)$, and if $R_{\Phi}$ is defined by the identity
$(\mathscr{L}\Phi_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}\Phi(x)|^{2}+\delta^{3}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+\delta^{3}R_{\Phi}(\delta;x),$
then
$\lim_{\delta\to 0}\sup\left\{|R_{\Phi}(\delta;x)|\,\mid\,x\in\mathbb{R}\right\}=0.$
The proof of the proposition is relegated to the end of this section.
There is one last technicality to address, namely, whereas Theorem 3 asks for the continuity of the initial datum $F_{\text{in}}$, the main theorem of this section allows for discontinuities. This gap is bridged using the fact that the equations of interest here are regularizing. This is made precise in the next condition:
  • (v)
    Asymptotically regularizing: If $\underline{F}$ and $\overline{F}$ denote the minimal and maximal viscosity sub- and supersolutions of the initial-value problem (52) for a given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ (see Appendix A.1), then
    $\underline{F}=\overline{F}\quad\text{in}\,\,\mathbb{R}\times(0,\infty).$
The next corollary asserts that if the family $\{T_{\epsilon}\}$ satisfies assumption (v) in addition to (i)-(iv), then Theorem 3 remains true even if $F_{\text{in}}$ has discontinuities.
@@
Assume that $\{T_{\epsilon}\}$ is a family of operators on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(v) above. Then the statement of Theorem 3 remains true for an arbitrary $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ provided $F$ is taken to be the unique discontinuous viscosity solution of the initial-value problem (52).
@@
In Theorem 2, the nonlinearity $\mathcal{F}$ takes the form (48), which satisfies assumption (v) by Propositions 10 and 11 below. Therefore, the proof of the theorem is complete.
@@
Theorem 3 is stated for semigroups defined in $CDF(\overline{\mathbb{R}})$, the space of nondecreasing, right-continuous functions taking values in $[0,1]$. There is nothing special about right-continuity, however. In fact, let $F_{-}$ and $F^{+}$ denote the left- and right-increasing versions of a nondecreasing function $F$. If operators $\{T_{\epsilon,-}\}$ are defined on the space $CDF_{\text{left}}(\overline{\mathbb{R}})$ of nondecreasing, left-continuous functions taking values in $[0,1]$ by the formula $T_{\epsilon,-}F=(T_{\epsilon}(F^{+}))_{-}$, then $\{T_{\epsilon,-}\}$ satisfies assumptions (i)-(iv) if and only if $\{T_{\epsilon}\}$ does. Further, it is not hard to see that a PDE scaling limit of $\{T_{\epsilon}\}$ holds as in Theorem 3 if and only if the same conclusion applies to $\{T_{\epsilon,-}\}$.

### v2

@@ 4.3. $\mathbb{Z}$-valued Solutions and Finite-Difference Equations
where $g_{\mathbb{Z}}$ denotes the left-inverse of the function $u\mapsto u+f_{\mathbb{Z}}(u)$. Thus, such a law $\mathbf{P}$ satisfies the assumptions of Theorems 1 and 2. In particular, as long as $(f_{+},f_{-})$ are not identically zero, either $n^{-\frac{1}{2}}X^{(n)}$ or $n^{-\frac{1}{3}}X^{(n)}$ converges to a nondegenerate limit.
@@ 4.4. Hipster Random Walks and Cooperative Motion
In fact, by this reasoning, it is also possible to draw a connection with the other RDE considered in [1], called the totally asymmetric hipster random walk. This is the RDE in $\mathbb{Z}$ with the one-step update determined by two independent Bernoulli random variables $(\Theta,\Xi)$ via the rule
@@
In view of the proposition above, the $\text{Beta}(2,2)$ limit theorems for the symmetric RDE’s considered in [1] and [3] follow as a special case of this paper’s main results. The $\text{Beta}(2,1)$ limit theorems for asymmetric RDE’s obtained in [1] and [2] also follow, albeit only in the parameter regimes identified above. Interestingly, in the asymmetric case, those works were able to obtain limit theorems even in the nonmonotone regime.
@@ 5. Scaling Limit
This section proves the main results, Theorems 1 and 2, as well as Corollary 1.
Fix data $(f_{+},f_{-})$ satisfying the assumptions of Theorems 1 or 2. Given $\theta\in\mathbb{R}$, define a sequence $\{p^{(N)}\}$ by
$p^{(N)}=1/2+\theta N^{-1}.$
Let $T^{(N)}$ be the operator associated with the RDE (9) with $\Phi$ given by (8) and bias parameter $p=p^{(N)}$. In view of the results of Section 3, the equation $F_{n}=T^{(N)}F_{n-1}$ can be written in the form
(47) $F_{n}-F_{n-1}=\mathcal{L}^{(N)}F_{n-1}-2\theta N^{-1}F_{n-1}(1-F_{n-1})$
for some function $\mathcal{L}^{(N)}$ defined on $CDF(\overline{\mathbb{R}})$. This equation, together with the properties of $\mathcal{L}^{(N)}$ obtained in Section 3, motivates the study of scaling limits of certain sequences of monotone operators defined on $CDF(\overline{\mathbb{R}})$ in an abstract setting.
5.1. Abstract Framework
Let $\{T^{(N)}\}$ be a family of operators on $CDF(\overline{\mathbb{R}})$ with the following properties:
  • (i)
    Monotonicity: If $F\leq G$ pointwise in $\mathbb{R}$, then $T^{(N)}F\leq T^{(N)}G$ also holds.
  • (ii)
    Discrete-Time Reaction-Diffusion Equation: There are functions $\{\mathcal{L}^{(N)}\}$ defined on $CDF(\overline{\mathbb{R}})$ and functions $\{Q^{(N)}\}$ defined on $[0,1]$ such that
    $T^{(N)}F-F=\mathcal{L}^{(N)}F-Q^{(N)}(F).$
  • (iii)
    Commutation with Constants: Given $F,G\in CDF(\overline{\mathbb{R}})$, if the difference FGF-G is a constant function, then $\mathcal{L}^{(N)}F=\mathcal{L}^{(N)}G$.
  • (iv)
    Action at Infinity: If $F\in CDF(\overline{\mathbb{R}})$, then, for each $\bar{x}\in\{+\infty,-\infty\}$,
    $(T^{(N)}F)(\bar{x})=F(\bar{x})-Q^{(N)}(F(\bar{x})).$
Notice that, as in the proof of Proposition 7 above, assumption (iv) implies that $\mathcal{L}^{(N)}$ vanishes on constant functions and, thus, $T^{(N)}$ maps constant functions to constant functions via the dynamical system
$T^{(N)}q=q-Q^{(N)}(q)\quad\text{for each}\,\,q\in[0,1].$
Since $T^{(N)}$ is monotone and maps $CDF(\overline{\mathbb{R}})$ to itself, it follows that the function $q\mapsto q-Q^{(N)}(q)$ is a nondecreasing map that sends $[0,1]$ into itself.
In addition to (i)-(iv) above, assume that there is a sequence of positive numbers $\{\delta^{(N)}\}$ such that $\delta^{(N)}\to 0$ as $N\to+\infty$ and a continuous function $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, which is nondecreasing in the second variable, for which the following property holds:
  • (v)
    Consistency of $\mathcal{L}^{(N)}$: If $\Phi\in CDF(\overline{\mathbb{R}})$ is smooth with bounded, uniformly continuous second derivatives, rescaled according to the formula $\Phi_{\delta}(y)=\Phi(\delta y)$, then
    $\lim_{N\to+\infty}N(\mathscr{L}^{(N)}\Phi_{\delta^{(N)}})\left(\frac{x}{\delta^{(N)}}\right)=\mathcal{F}(\partial_{x}\Phi(x),\partial_{x}^{2}\Phi(x)),$
    where convergence holds uniformly with respect to $x\in\mathbb{R}$.
Since $\mathcal{L}^{(N)}$ vanishes on constant functions for any $N$, assumption (v) implies
(48) $\mathcal{F}(0,0)=0.$
An analogous assumption will be made on the reaction term $Q^{(N)}$, as follows:
  • (vi)
    Consistency of $Q^{(N)}$: There is a function $Q:[0,1]\to[0,1]$ such that
    $\lim_{N\to\infty}\sup\left\{|NQ^{(N)}(q)-Q(q)|\,\mid\,q\in[0,1]\right\}=0$
    Further, there is a constant $L>0$ such that, independently of $N$,
    $|Q^{(N)}(q)-Q^{(N)}(q^{\prime})|\leq LN^{-1}|q-q^{\prime}|\quad\text{for each}\,\,q,q^{\prime}\in[0,1].$
Under these assumptions, sequences of solutions $\{F^{(N)}_{n}\}$ of the recursion $F^{(N)}_{n}=T^{(N)}F^{(N)}_{n-1}$ undergo a scaling limit, whereby their asymptotic behavior is determined by the PDE $\partial_{t}F=\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)-Q(F)$. This is made precise in the next theorem.
@@ 5.1. Abstract Framework
Assume that $\{T^{(N)}\}$ is a family of operators on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(vi) above. For each $N$, let $\{F^{(N)}_{n}\}$ be a sequence in $CDF(\overline{\mathbb{R}})$ generated by the recursion $F_{n}^{(N)}=T^{(N)}F^{(N)}_{n-1}$. If $F_{N}:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is the rescaled CDF defined by
$F_{N}(x,t)=F^{(N)}_{[Nt]}\left(\frac{x}{\delta^{(N)}}\right),$
and if there is a continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ such that, at time $t=0$,
$F^{(N)}(\cdot,0)\to F_{\text{in}}\quad\text{vaguely as}\,\,N\to+\infty,$
@@
$F_{N}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times[0,\infty)\,\,\text{as}\,\,N\to+\infty,$
where $F$ is the unique bounded viscosity solution of the equation
(49) $\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
It is possible to show that if $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, as is assumed in Theorems 1 and 2, then the one-parameter family $\{T^{(N)}\}$ defined in the discussion at the beginning of this section satisfies assumptions (i)-(iv); see Section 5.2 for the details. In particular, $\{Q^{(N)}\}$ is given by $Q^{(N)}(q)=2\theta N^{-1}q(1-q)$, which trivially satisfies (vi). It remains to verify condition (v). That is the subject of the next proposition. Recall that the constants $\sigma$ and $a$ are defined in Section 1.4.1 above.
@@
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, and fix $\theta\in\mathbb{R}$. Define $p^{(N)}=1/2+\theta N^{-1}$. Consider the operator $\mathscr{L}^{(N)}$ given by (32) with $p=p^{(N)}$. If $\Phi:\mathbb{R}\to[0,1]$ is a nondecreasing function with bounded, uniformly continuous first and second derivatives, rescaled according to the rule $\Phi_{\delta}(y)=\Phi(\delta y)$, and if $R_{\Phi}^{(N)}:\mathbb{R}\to\mathbb{R}$ is the remainder defined by
$(\mathscr{L}^{(N)}\Phi_{N^{-\frac{1}{2}}})(N^{\frac{1}{2}}x)=N^{-1}\sigma|\partial_{x}\Phi(x)|^{2}+N^{-1}R_{\Phi}^{(N)}(x),$
then $R^{(N)}_{\Phi}\to 0$ uniformly in $\mathbb{R}$ as $N\to+\infty$.
Further, if the stronger assumption (16) holds, and if $R^{(N)}_{\Phi}$ is instead defined by
$(\mathscr{L}^{(N)}\Phi_{N^{-\frac{1}{3}}})(N^{\frac{1}{3}}x)=N^{-\frac{2}{3}}\sigma|\partial_{x}\Phi(x)|^{2}+N^{-1}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+N^{-1}R_{\Phi}^{(N)}(x),$
then, as before, $R^{(N)}_{\Phi}\to 0$ uniformly as $N\to+\infty$.
The proof of the proposition is relegated to the end of this section (Section 5.4).
There is one last technicality to address, namely, whereas Theorem 3 asks for the continuity of the initial datum $F_{\text{in}}$, Theorems 1 and 2 allow for discontinuities. This gap is bridged using the fact that the equations of interest here are well-posed even for discontinuous data. This is made precise in the next condition:
  • (vii)
    Asymptotically regularizing: If $\overline{F}$ and $\underline{F}$ denote the maximal and minimal viscosity sub- and supersolutions, respectively, of the initial-value problem (49) for a given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ (see Appendix A.1), then
    $\overline{F}=\underline{F}\quad\text{in}\,\,\mathbb{R}\times(0,\infty).$
As explained in Appendix A, from the fact that $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$, it follows that there is a unique solution of (49) started from any initial datum $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, which is continuous for positive times.
The next corollary asserts that if the family $\{T^{(N)}\}$ satisfies assumption (vii) in addition to (i)-(vi), then Theorem 3 remains true even if $F_{\text{in}}$ has discontinuities.
@@
Assume that $\{T^{(N)}\}$ is a family of operators on $CDF(\overline{\mathbb{R}})$ satisfying assumptions (i)-(vii) above. Then the statement of Theorem 3 remains true for an arbitrary $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ provided $F$ is taken to be the unique bounded discontinuous viscosity solution of the initial-value problem (49).
@@
It follows from classical results that, in the PDE’s of interest here, assumption (vii) holds as long as either $\sigma\neq 0$ or $a>0$. (The degenerate PDE’s $\partial_{t}F-Q(F)=0$ do not have this property and therefore require a slightly different treatment.) This is made precise in the next proposition.
Proposition 11.
Let $Q$ be an arbitrary function for which there is an $L>0$ such that $|Q(q)-Q(q^{\prime})|\leq L|q-q^{\prime}|$ for each $q,q^{\prime}\in[0,1]$ and satisfying both $Q(0)\geq 0$ and $Q(1)\leq 0$. If $\mathcal{F}(v,w)=\sigma|v|^{2}$ for some $\sigma\neq 0$ or $\mathcal{F}(v,w)=a|v|w$ for some $a>0$, then assumption (vii) is satisfied.
Proof.
When $\theta=0$, this is established in Propositions 14 and 15 in Appendix A. The improvement to the case $\theta\neq 0$ is covered by Proposition 16. ∎
@@
Theorem 3 is stated for semigroups defined in $CDF(\overline{\mathbb{R}})$, the space of nondecreasing, right-continuous functions taking values in $[0,1]$. There is nothing special about right-continuity, however. In fact, let $F_{-}$ and $F^{+}$ denote the left- and right-increasing versions of a nondecreasing function $F$. If operators $\{T^{(N)}_{-}\}$ are defined on the space $CDF_{\text{left}}(\overline{\mathbb{R}})$ of nondecreasing, left-continuous functions taking values in $[0,1]$ by the formula $T^{(N)}_{-}F=(T^{(N)}(F^{+}))_{-}$, then $\{T^{(N)}_{-}\}$ satisfies assumptions (i)-(vi) if and only if $\{T^{(N)}\}$ does. Further, it is not hard to see that a PDE scaling limit of $\{T^{(N)}\}$ holds as in Theorem 3 if and only if the same conclusion applies to $\{T^{(N)}_{-}\}$.
5.2. Proof of Theorems 1 and 2
Both theorems are direct applications of Theorem 3, or rather its corollary, Corollary 4.
As in the discussion at the beginning of this section, let $T^{(N)}$ be the operator associated with the RDE (9) with data $(f_{+},f_{-})$ satisfying $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) and bias parameter $p^{(N)}=1/2+\theta N^{-1}$ for some fixed constant $\theta\in\mathbb{R}$. Assumptions (i) and (ii) follow from Propositions 2 and 5, respectively.
The assumptions (iii) and (iv) hold by Proposition 7. To see that this proposition is applicable, it only remains to check that $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. Fix $\rho\in\{+,-\}$. Since $s-g_{\rho}(s)=f_{\rho}(g_{\rho}(s))$ for $s\geq f_{\rho}(0)$ and $g_{\rho}(s)\to+\infty$ as $s\to+\infty$, the integral $\int_{f_{\rho}(0)}^{+\infty}|g_{\rho}(s)-s|\,ds$ is finite only if $f_{\rho}(+\infty)=0$. At the same time, (15) implies that this integral is finite almost surely, hence $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$ holds as claimed. Thus, the hypotheses of Proposition 7 are satisfied, and it follows that assumptions (iii) and (iv) both apply to the sequence $\{T^{(N)}\}$.
In the context of Theorem 1 (i.e., assuming both (15) and $\sigma\neq 0$), Proposition 10 implies that assumption (v) holds with
$\delta^{(N)}=N^{-\frac{1}{2}}\quad\text{and}\quad\mathcal{F}(v,w)=\sigma|v|^{2}.$
According to Proposition 11 above, from the fact that $\sigma\neq 0$, assumption (vii) also holds. Therefore, Corollary 4 applies, establishing convergence of the rescaled CDF $F_{N}$ as desired to the solution of the initial-value problem associated to the PDE $\partial_{t}F-\sigma|\partial_{x}F|^{2}+2\theta F(1-F)=0$, as desired.
Similarly, in Theorem 2 (that is, assuming that (16) holds, $\sigma=0$, and $a>0$), Proposition 10 instead establishes that assumption (v) holds with
$\delta^{(N)}=N^{-\frac{1}{3}}\quad\text{and}\quad\mathcal{F}(v,w)=a|v|w.$
Once again, because $a>0$, Proposition 11 asserts that assumption (vii) is satisfied. Therefore, by Corollary 4, the rescaled CDF $F_{N}$ instead converges to the solution of the initial-value problem associated to $\partial_{t}F-a|\partial_{x}F|\partial_{x}^{2}F+2\theta F(1-F)=0$. ∎